# PipeNet: Question Answering with Semantic Pruning over Knowledge Graphs

### v1

@@ Abstract
Ying Su, Jipeng Zhang, Yangqiu Song, Tong Zhang HKUST {ysuay,jzhanggr}@connect.ust.hk, yqsong@cse.ust.hk, tongzhang@ust.hk
@@ 1 Introduction
Experiments on two standard benchmarks, CommonsenseQA Talmor et al. (2019) and OpenbookQA Mihaylov et al. (2018), demonstrate the effectiveness of our proposed method.
@@ 2.1 QA with LM+KG
With the development of benchmarking question answering, more and more hard question answering datasets are developed, which require background knowledge to solve Mihaylov et al. (2018); Talmor et al. (2019, 2021). Pretrained LMs and KGs are commonly used knowledge sources, and recent research mainly adopts an LM+KG framework for encoding the query context and KG nodes.
@@ 3.2 DP-pruning
     if $\mathcal{D}_{t}[v]\geq threshold$ then
@@
Propogate node score. Our algorithm aims to prune the external nodes $\mathcal{V}_{e}$ in the subgraph for two-hop or above because noisy nodes are mainly induced with the hops growing. The pseudo-code for pruning the external nodes is listed in Algorithm 1. After initialization with grounded concept sets $\mathcal{V}_{t}=\mathcal{V}_{q}\bigcup\mathcal{V}_{a}$, as well as node score set for grounded concept sets $\mathcal{D}_{t}=\mathcal{D}_{q}\bigcup\mathcal{D}_{a}$. External nodes having neighbors in the grounded concept sets are added to expand the grounded subgraph $\mathcal{G}_{sub}$. The node score of external nodes is assigned as the average of their neighbor node scores during expansion. Until the expansion hops end, the nodes except $\mathcal{V}_{q}$ and $\mathcal{V}_{a}$ are pruned according to their score ranks. The nodes with smaller node scores are pruned.
@@ 3.3 Reasoning
$L$-h KagNet
$\mathcal{O}(|\mathcal{R}|^{L}|\mathcal{V}|^{L+1}L)$
$\mathcal{O}(|\mathcal{R}|^{L}|\mathcal{V}|^{L+1}L\cdot D)$
$L$-h MHGRN
$\mathcal{O}(|\mathcal{R}|^{2}|\mathcal{V}|^{2}L)$
$\mathcal{O}(|\mathcal{R}||\mathcal{V}|L\cdot D)$
$L$-l QAGNN
$\mathcal{O}(|\mathcal{V}|^{2}L)$
$\mathcal{O}(|\mathcal{R}||\mathcal{V}|L\cdot D)$
$L$-l GreaseLM
$\mathcal{O}(|\mathcal{V}|^{2}L)$
$\mathcal{O}(|\mathcal{R}||\mathcal{V}|L\cdot D)$
$L$-l JointLK
$\mathcal{O}(|\mathcal{V}|^{2}L)$
$\mathcal{O}(|\mathcal{R}||\mathcal{V}|L\cdot D)$
$L$-l GSC
$\mathcal{O}(|\mathcal{V}|L)$
$\mathcal{O}(|\mathcal{R}||\mathcal{V}|L)$
$L$-l PipeNet
$\mathcal{O}((\frac{|\mathcal{V}|}{k})^{2}L)$
$\mathcal{O}(|\mathcal{R}|\frac{|\mathcal{V}|}{k}L\cdot D)$
@@
$L$-h KagNet
$\mathcal{O}(|\mathcal{R}|^{L}|\mathcal{V}|L\Delta^{L})$
$\mathcal{O}(|\mathcal{R}|^{L}||\mathcal{V}|L\Delta^{L}\cdot D)$
$L$-h MHGRN
$\mathcal{O}(|\mathcal{R}|^{2}|\mathcal{V}|L\Delta)$
$\mathcal{O}(|\mathcal{R}||\mathcal{V}|L\cdot D)$
$L$-l QAGNN
$\mathcal{O}(|\mathcal{V}|L\Delta)$
$\mathcal{O}(|\mathcal{R}||\mathcal{V}|L\cdot D)$
$L$-l GreaseLM
$\mathcal{O}(|\mathcal{V}|L\Delta)$
$\mathcal{O}(|\mathcal{R}||\mathcal{V}|L\cdot D)$
$L$-l JointLK
$\mathcal{O}(|\mathcal{V}|L\Delta)$
$\mathcal{O}(|\mathcal{R}||\mathcal{V}|L\cdot D)$
$L$-l GSC
$\mathcal{O}(|\mathcal{V}|L)$
$\mathcal{O}(|\mathcal{R}||\mathcal{V}|L)$
$L$-l PipeNet
$\mathcal{O}(\frac{|\mathcal{V}|}{k}L\Delta)$
$\mathcal{O}(|\mathcal{R}|\frac{|\mathcal{V}|}{k}L\cdot D)$
@@
$z$ is the hidden state of [CLS] token in the last hidden layer.
@@ 4.3 Implementation Details
For all the experiments on PipeNet, we set the dimension ($D=200$) and the number of layers ($L=5$) in the reasoning module. The parameters of the reasoning module (LM+GNN) are optimized by RAdam Liu et al. (2019a) by cross-entropy loss. The learning rate for the LM encoder is set as 1e-5. For the decoder with GNN, the learning rate is 1e-3. For both benchmarks, we use ConceptNet Speer et al. (2017) as the knowledge graph. For the pruning experiments on PipeNet, the DP-pruning strategy prunes the nodes by the ranks of node scores. For each experiment, we use two RTX 3090 GPUs, and the average running time is about 4 hours on CSQA and 24 hours on OBQA.
@@ 5.2 Efficiency of PipeNet and DP-pruning
$p$(%)
$k$ $M$(G) $\uparrow$(%) $t$(s) $\uparrow$(%) IHtest (%)
0
1.0 5.02 - 1.16 - 74.38
10
1.1 4.95 1 1.02 13 74.21
20
1.2 4.92 2 1.01 13 74.29
30
1.3 4.75 5 0.96 17 74.29
40
1.4 4.67 7 0.87 25 74.21
50
1.6 4.57 9 0.83 28 74.13
60
2.0 4.49 11 0.78 33 74.70
70
2.4 4.22 16 0.75 35 74.85
80
3.2 3.83 24 0.72 38 74.70
90
4.8 3.51 30 0.67 42 74.86
@@ 5.3 Ablation Study
$h$-hop
Prune method
Prune rate
IHtest-Acc(%)
One
0
73.27 (±$\pm 0.93$)
Two
0
74.49 (±$\pm 0.26$)
Two
Random
90% 73.51 (±$\pm 0.61$)
Two
DP
90% 74.75 (±$\pm 0.47$)

### v2

@@ Abstract
Ying Su, Jipeng Zhang, Yangqiu Song, Tong Zhang Hong Kong University of Science and Technology {ysuay,jzhanggr}@connect.ust.hk, yqsong@cse.ust.hk, tongzhang@ust.hk
@@ 1 Introduction
Experiments on two standard benchmarks, CommonsenseQA Talmor et al. (2019) and OpenbookQA Mihaylov et al. (2018), demonstrate the effectiveness of our proposed method. The code is open-sourced [^1] https://github.com/HKUST-KnowComp/PipeNet.
@@ 2.1 QA with LM+KG
With the development of benchmarking question answering, more and more hard question answering datasets are developed, which require background knowledge to solve Mihaylov et al. (2018); Talmor et al. (2019, 2021). Pretrained LMs and KGs are commonly used knowledge sources, research typically adopts an LM+KG framework as to acquire relevent knowledge for commonsense QA Feng et al. (2020); Yasunaga et al. (2021); Zhang et al. (2022b); Su et al. (2022); Park et al. (2023); Huang et al. (2023); Ye et al. (2023); Wang et al. (2023); Taunk et al. (2023); Zhao et al. (2023); Dong et al. (2023); Mazumder and Liu ; Kang et al. (2024); Zhao et al. (2024)
@@ 3.2 DP-pruning
     if $\mathcal{D}_{t}[v]\leq threshold$ then
@@
Propogate node score. Our algorithm aims to prune the external nodes $\mathcal{V}_{e}$ in the subgraph for two-hop or above because noisy nodes are mainly induced with the hops growing. The pseudo-code for pruning the external nodes is listed in Algorithm 1. In initialization, grounded concept sets $\mathcal{V}_{t}=\mathcal{V}_{q}\bigcup\mathcal{V}_{a}$, and score set for grounded concept sets $\mathcal{D}_{t}=\mathcal{D}_{q}\bigcup\mathcal{D}_{a}$. External nodes having neighbors in the grounded concept sets are added to expand the grounded subgraph $\mathcal{G}_{sub}$. The node score of external nodes is assigned as the average of their neighbor node scores during expansion. Until the expansion hops end, the nodes except $\mathcal{V}_{q}$ and $\mathcal{V}_{a}$ are pruned according to their score ranks. The nodes with smaller node scores are pruned.
@@ 3.3 Reasoning
$L$-h KagNet $\mathcal{O}(|\mathcal{R}|^{L}|\mathcal{V}|^{L+1}L)$ $\mathcal{O}(|\mathcal{R}|^{L}|\mathcal{V}|^{L+1}L\cdot D)$
$L$-h MHGRN $\mathcal{O}(|\mathcal{R}|^{2}|\mathcal{V}|^{2}L)$ $\mathcal{O}(|\mathcal{R}||\mathcal{V}|L\cdot D)$
$L$-l QAGNN $\mathcal{O}(|\mathcal{V}|^{2}L)$ $\mathcal{O}(|\mathcal{R}||\mathcal{V}|L\cdot D)$
$L$-l GreaseLM $\mathcal{O}(|\mathcal{V}|^{2}L)$ $\mathcal{O}(|\mathcal{R}||\mathcal{V}|L\cdot D)$
$L$-l JointLK $\mathcal{O}(|\mathcal{V}|^{2}L)$ $\mathcal{O}(|\mathcal{R}||\mathcal{V}|L\cdot D)$
$L$-l GSC $\mathcal{O}(|\mathcal{V}|L)$ $\mathcal{O}(|\mathcal{R}||\mathcal{V}|L)$
$L$-l PipeNet $\mathcal{O}((\frac{|\mathcal{V}|}{k})^{2}L)$ $\mathcal{O}(|\mathcal{R}|\frac{|\mathcal{V}|}{k}L\cdot D)$
@@
$L$-h KagNet $\mathcal{O}(|\mathcal{R}|^{L}|\mathcal{V}|L\Delta^{L})$ $\mathcal{O}(|\mathcal{R}|^{L}||\mathcal{V}|L\Delta^{L}\cdot D)$
$L$-h MHGRN $\mathcal{O}(|\mathcal{R}|^{2}|\mathcal{V}|L\Delta)$ $\mathcal{O}(|\mathcal{R}||\mathcal{V}|L\cdot D)$
$L$-l QAGNN $\mathcal{O}(|\mathcal{V}|L\Delta)$ $\mathcal{O}(|\mathcal{R}||\mathcal{V}|L\cdot D)$
$L$-l GreaseLM $\mathcal{O}(|\mathcal{V}|L\Delta)$ $\mathcal{O}(|\mathcal{R}||\mathcal{V}|L\cdot D)$
$L$-l JointLK $\mathcal{O}(|\mathcal{V}|L\Delta)$ $\mathcal{O}(|\mathcal{R}||\mathcal{V}|L\cdot D)$
$L$-l GSC $\mathcal{O}(|\mathcal{V}|L)$ $\mathcal{O}(|\mathcal{R}||\mathcal{V}|L)$
$L$-l PipeNet $\mathcal{O}(\frac{|\mathcal{V}|}{k}L\Delta)$ $\mathcal{O}(|\mathcal{R}|\frac{|\mathcal{V}|}{k}L\cdot D)$
@@
where $z$ is the hidden state of [CLS] token in the last hidden layer.
@@ 4.3 Implementation Details
For all the experiments on PipeNet, we set the dimension ($D=200$) and the number of layers ($L=5$) in the reasoning module. The parameters of the reasoning module (LM+GNN) are optimized by RAdam Liu et al. (2019a) by cross-entropy loss. The learning rate for the LM encoder is set as 1e-5. For the decoder with GNN, the learning rate is 1e-3. For both benchmarks, we use ConceptNet Speer et al. (2017) as the knowledge graph. For the pruning experiments on PipeNet, the DP-pruning strategy prunes the nodes by the ranks of node scores. Specifically, the $threshold$ value is determined by the score of top $(1-p)$ percent ranked node is $\mathcal{V}_{t}\setminus\{\mathcal{V}_{q},\mathcal{V}_{a}\}$. In each experiment, we use two RTX 3090 GPUs, and the average running time is about 4 hours on CSQA and 24 hours on OBQA.
@@ 5.2 Efficiency of PipeNet and DP-pruning
$p$(%) $k$ $M$(G) $\uparrow$(%) $t$(s) $\uparrow$(%) IHtest (%)
0 1.0 5.02 - 1.16 - 74.38
10 1.1 4.95 1 1.02 13 74.21
20 1.2 4.92 2 1.01 13 74.29
30 1.3 4.75 5 0.96 17 74.29
40 1.4 4.67 7 0.87 25 74.21
50 1.6 4.57 9 0.83 28 74.13
60 2.0 4.49 11 0.78 33 74.70
70 2.4 4.22 16 0.75 35 74.85
80 3.2 3.83 24 0.72 38 74.70
90 4.8 3.51 30 0.67 42 74.86
@@ 5.3 Ablation Study
$h$-hop Prune method Prune rate IHtest-Acc(%)
One - 0 73.27 (±$\pm 0.93$)
Two - 0 74.49 (±$\pm 0.26$)
Two Random 90% 73.51 (±$\pm 0.61$)
Two DP 90% 74.75 (±$\pm 0.47$)
@@ 6 Conclusion
7 Acknowledgement
The authors of this paper were supported by the NSFC Fund (U20B2053) from the NSFC of China, the RIF (R6020-19 and R6021-20) and the GRF (16211520 and 16205322) from RGC of Hong Kong. We also thank the UGC Research Matching Grants (RMGS20EG01-D, RMGS20CR11, RMGS20CR12, RMGS20EG19, RMGS20EG21, RMGS23CR05, RMGS23EG08). We would like to thank the Turing AI Computing Cloud (TACC) Xu et al. (2021) and HKUST iSING Lab for providing us computation resources on their platform.